OpenStudy (anonymous):

given: 1 pail = 4 liters 1 pail = 9 liters You are going to the river to fetch water. How could you carry 6 liters of water using only this two pails? note: pails have no permanent calibration.

OpenStudy (mathmate):

Think of the equation 9x+4y=1 Look for x and y to satisfy the above equation, obviously one of them has to be negative. This will give you an idea how to go about getting the answer. If you need a systematic way to solve it, look up diophantine equations.

OpenStudy (anonymous):

pail 1 pail 2 9 0 5 4 5 0 1 4 1 0 0 1 9 1 6 4

OpenStudy (mathmate):

Sorry, the above equation should have read: 9x+4y=6 which makes 9*(2)-4*(3)=6 (using Gow's results)

OpenStudy (anonymous):

my solution is: (1) the 2 pails have no calibration (2) solve: 9 -4 =5 5-4 = 1 (3) now that you have remaining 1 liter of water. (4) put that 1 liter of water to 4 liter-capacity pail (5) the 9 liter-capacity pail has no content (6) full it with water (7) now, since the 4 liter-capacity pail has 1 liter inside, you need only 3 liter to full it. (8) solve: 9-3 = 6 ; final answer